# Human or LLM as Standardized Patients? A Comparative Study for Medical Education

### v1

@@ E.2.1 Part 1: Background Information
E.2.2 Evaluation of Learning and Training Models
@@
E.2.3 Overall Assessment and Open-ended Feedback
@@
Appendix F Experimental Settings
F.1 A. EasyMED (ours)
@@ Appendix F Experimental Settings
Appendix G User Interface of the EasyMED
@@
Appendix H Validation of the Feedback Agent’s Guidance
H.1 Objective
The feedback agent in our training platform provides post-session guidance to users, such as highlighting missed or redundant questions. To ensure that this guidance is appropriate and clinically sound, we conducted an independent validation study.
H.2 Study Design
  • Sample selection: We collected 30 complete practice sessions from participants using the platform.
  • Feedback extraction: For each session, we extracted the feedback agent’s generated guidance, including “missed items” and “superfluous items.”
  • Expert review: Two clinical experts, blinded to the feedback agent, independently rated the appropriateness of each feedback item on a 5-point Likert scale (1=misleading, 5=highly appropriate).
H.3 Evaluation Metrics
We computed:
  • Accuracy of feedback: Percentage of feedback items judged "appropriate" (score $\geq$ 4) by both experts.
  • Inter-rater agreement: Cohen’s $\kappa$ between the two experts for appropriateness labels.
  • Error categorization: Manual classification of disagreements (false positives, omissions, ambiguity).
H.4 Results Summary
Across 30 sessions, 87% of the feedback agent’s suggestions were rated appropriate (score $\geq$ 4 by both experts). Inter-rater agreement between the two experts was $\kappa=0.76$. Most disagreements arose in borderline cases where the clinical relevance of a missed question was debatable.
H.5 Discussion and Recommendations
These findings indicate that the feedback agent provides generally reliable guidance but may require refinement for ambiguous or edge cases. Future work could incorporate a confidence score or allow experts to flag uncertain feedback for model retraining. All annotated feedback and rating sheets are included in the supplementary materials for replication.
Appendix I Core Prompts Used in the Study
@@
I.1 Patient Agent Prompt
@@
I.2 Auxiliary Agent Prompt
@@
I.3 Evaluation Agent Prompt
@@
I.4 Automated Evaluation Prompt

### v2

@@ F.2.1 Part 1: Background Information
F.2.2 Evaluation of Learning and Training Models
@@
F.2.3 Overall Assessment and Open-ended Feedback
@@
Appendix G Participant Exclusion
We initially recruited 20 medical students. Before random group assignment and prior to any training, six participants were excluded based on predefined criteria, resulting in a final sample of 14 students.
Specifically, four students were excluded due to scheduling conflicts that prevented them from attending the required in-person human SP sessions, and two students were excluded due to extreme pre-test OSCE scores (95 and 96 out of 100), where ceiling effects would limit measurable learning gains. All exclusions occurred before group assignment and independently of the intervention, ensuring no differential attrition between conditions. Although the excluded students had a higher mean baseline score than the included cohort, this does not introduce selection bias because exclusions were applied prior to randomization.
Appendix H Experimental Settings
H.1 A. EasyMED (ours)
@@ Appendix H Experimental Settings
Appendix I User Interface of the EasyMED
@@
Appendix J Core Prompts Used in the Study
@@
J.1 Automated Evaluation with GPT-4o
To enable scalable and reproducible evaluation on SPBench, we employ GPT-4o as an automated judge to assess the quality of virtual standardized patient responses. This section details the evaluation pipeline, including model inputs, scoring procedure, and score aggregation.
Evaluation Input.
For each test instance, GPT-4o is provided with three components: (1) a structured case describing the ground-truth patient profile, (2) the full doctor–patient dialogue transcript, and (3) a fixed evaluation prompt specifying eight expert-defined evaluation criteria. The same prompt and input format are used for all evaluated systems to ensure consistency.
Scoring Procedure.
GPT-4o evaluates the patient responses independently along eight dimensions (Query Comprehension, Case Consistency, Controlled Disclosure, Response Completeness, Logical Coherence, Language Naturalness, Conversational Consistency, and Patient Demeanor). Each dimension is rated on a 5-point Likert scale following explicit rubric definitions. The model is instructed to justify each score by citing specific dialogue turns as evidence and to return the results in a structured JSON format.
Score Aggregation.
For reporting, raw Likert scores are linearly rescaled to a 0–100 scale. Dimension-level scores are averaged across all test dialogues, and an overall score is computed as the mean of the eight dimension scores. No manual intervention or post-hoc adjustment is applied during this process.
Automated Professional A Professional B
Average Score 84.1 91.3 86.4
Standard Deviation 8.5 9.1 9.3
Correlation 0.81
Table 10: Inter-rater reliability between the automated evaluator (GPT-4o) and human clinical experts, measured by Pearson correlation
LLM vs. Human Evaluation
Considering the high cost of large-scale manual annotation, we used GPT-4o as an automated judge. We gave it a carefully written prompt (Appendix J.5). To verify its reliability, we randomly selected 86 samples for blind review by two clinical experts (see Appendix  LABEL:sec:_partner for rater qualifications) who were unaware of the GPT-4o ratings. The automated scores correlated strongly and significantly with the experts’ average ratings (Pearson’s $r=0.81$; Table 10). Although GPT-4o’s mean score (84.1) was consistently lower than the experts’ mean scores (88.9), this high alignment indicates that GPT-4o provides a reliable proxy for large-scale evaluation.
J.2 Patient Agent Prompt
@@
J.3 Auxiliary Agent Prompt
@@
J.4 Evaluation Agent Prompt
@@
J.5 Automated Evaluation Prompt